given x= 0.8531 population mean= 0.8596 standard deviation = 0.0522 n= 460 find the z score...

given
x= 0.8531
population mean= 0.8596
standard deviation = 0.0522
n= 460

find the z score and the accumulative area
Answer #1

z score for x = 0.8531:

z = (0.8531 - 0.8596)/(0.0522/) = -2.67

Cumulative area = P(z < -2.67) = 0.0038